OpenStudy (anonymous):

will medal For f(x) = 0.01(2)x, find the average rate of change from x = 2 to x = 10. 1.275 8 10.2 10.24

OpenStudy (anonymous):

@Keigh2015 help quick

OpenStudy (anonymous):

also Which of the following options results in a graph that shows exponential growth? f(x) = 0.4(3)x f(x) = 3(0.5)x f(x) = 0.8(0.9)x f(x) = 0.9(5)−x

OpenStudy (anonymous):

@Keigh2015

OpenStudy (anonymous):

@BPDlkeme234

OpenStudy (anonymous):

\[\frac{ 0.2 }{ x }\] I think... This one is hard.

OpenStudy (anonymous):

Let's solve for y. y=(0.01)(2)(10) Answer: y=0.2

OpenStudy (anonymous):

do u know anyone else who can help

OpenStudy (anonymous):

srry

OpenStudy (anonymous):

go on

OpenStudy (anonymous):

@mathmate @misssunshinexxoxo @MICAAAAAAAAAAAAAAAAAA

OpenStudy (anonymous):

keep going

OpenStudy (anonymous):

@misssunshinexxoxo

OpenStudy (anonymous):

@Keigh2015

OpenStudy (anonymous):

I believe it is 10.2.... I'm not sure.

OpenStudy (keigh2015):

I agree with vbroberson

OpenStudy (anonymous):

Okay. I'm just not sure.

OpenStudy (anonymous):

this one too which of the following options results in a graph that shows exponential growth? f(x) = 0.4(3)x f(x) = 3(0.5)x f(x) = 0.8(0.9)x f(x) = 0.9(5)−x

OpenStudy (anonymous):

@Keigh2015

OpenStudy (keigh2015):

hmm let me think

OpenStudy (keigh2015):

I would say the last one

OpenStudy (anonymous):

u sure

OpenStudy (anonymous):

it was wrong :( but thanks anyway

OpenStudy (keigh2015):

I am sorry

OpenStudy (anonymous):

sokay

OpenStudy (anonymous):

try this one Which of the following options results in a graph that shows exponential growth? f(x) = 0.4(0.2)x f(x) = 4(0.98)x f(x) = 0.7(5)x f(x) = 5(4)−x

OpenStudy (keigh2015):

Not the last one I know that

OpenStudy (mathmate):

For \(f(x) = 0.01(2)^x\), find the average rate of change from x = 2 to x = 10. For this one, you can use average rate of change = (f(x2)-f(x1))/(x2-x1) The correct answer is one of the choices. Which of the following options results in a graph that shows exponential growth? f(x) = 0.4(0.2)x f(x) = 4(0.98)x f(x) = 0.7(5)x f(x) = 5(4)−x For this one,you need to be exact with the question. Whether the x is a power of a simple multiplication makes the whole difference. So please type the exact choices. (0.2)^x = 0.2 raised to the power of x. (0.2)x means 0.2 multiplied by x.